OpenStudy (anonymous):

Somebody seriously please help me. What is the excluded value of the rational expression (x - 4)/(6) ÷ (x + 1)/(7)?

OpenStudy (mertsj):

Any value that makes the denominator 0 is excluded.

OpenStudy (anonymous):

I know but I don't know what or how to find the excluded value of this rational expression @Mertsj

OpenStudy (mertsj):

As it is written, the denominators are constants and can never be 0. However, if you attempt to do the computation, the second fraction will have to be inverted and then the denominator will be x+1.

OpenStudy (anonymous):

Is this right? 7(x + 4)/ 6(x +1)

OpenStudy (mertsj):

You know that dividing is the same a multiplying by the reciprocal, right?

OpenStudy (anonymous):

(69+69)/ 6(9x+69)

OpenStudy (anonymous):

Yes I think

OpenStudy (anonymous):

? @SemiColonProject

OpenStudy (mertsj):

\[\frac{x-4}{6}\div \frac{x+1}{7}=\frac{x-4}{6}\times\frac{7}{x+1}\]

OpenStudy (mertsj):

Does that look familiar?

OpenStudy (anonymous):

A little

OpenStudy (mertsj):

So what number would cause the denominator to be 0?

OpenStudy (anonymous):

Can you show me the next step then I would know what number would cause the denominator to be 0.

OpenStudy (mertsj):

The denominator of concern is x+1. Do you know that?

OpenStudy (anonymous):

Okay

OpenStudy (mertsj):

Because the other denominator is 6 and 6 is never 0

OpenStudy (mertsj):

So talking about the denominator x+1...Would it be 0 if x is 0?

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

No

OpenStudy (mertsj):

Right. Very good.

OpenStudy (mertsj):

If we replace x with 1, will x+1 be 0?

OpenStudy (anonymous):

No

OpenStudy (mertsj):

So what number WILL make x+1 be 0?

OpenStudy (anonymous):

Here are the answer choices for this question: A. -1 B. 4 C. -7 D. None of the above.

OpenStudy (mertsj):

We could write x+1=0 and solve it to find out. Can you solve that?

OpenStudy (anonymous):

I got it!!!!!!!!! It's -1

OpenStudy (mertsj):

Very good. You are very smart.

OpenStudy (anonymous):

Thanks. So -1 is the answer right? Just to be sure.

OpenStudy (mertsj):

Yes it is.

OpenStudy (mertsj):

And you are welcome.

OpenStudy (anonymous):

I got the problem right. Thanks again.

OpenStudy (anonymous):

@Mertsj